# LLM-based speaker diarization correction: A generalizable approach

### v1

@@ Abstract
Speaker diarization is necessary for interpreting conversations transcribed using automated speech recognition (ASR) tools. Despite significant developments in diarization methods, diarization accuracy remains an issue. Here, we investigate the use of large language models (LLMs) for diarization correction as a post-processing step. LLMs were fine-tuned using the Fisher corpus, a large dataset of transcribed conversations. The ability of the models to improve diarization accuracy in a holdout dataset was measured. We report that fine-tuned LLMs can markedly improve diarization accuracy. However, model performance is constrained to transcripts produced using the same ASR tool as the transcripts used for fine-tuning, limiting generalizability. To address this constraint, an ensemble model was developed by combining weights from three separate models, each fine-tuned using transcripts from a different ASR tool. The ensemble model demonstrated better overall performance than each of the ASR-specific models, suggesting that a generalizable and ASR-agnostic approach may be achievable. We hope to make these models accessible through public-facing APIs for use by third-party applications.
Index Terms:
speaker diarization, error correction, large language model
Preprint Notice:
This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
I Introduction
@@
Here, we fine-tune LLMs to improve diarization accuracy in ASR transcripts as a post-processing step. To both fine-tune and test our models, we used the English Fisher corpus [31], a large audio dataset containing 1,960 hours of phone conversations. We consider the effects of using transcripts from different ASRs, tuning separate models for each AWS, Azure, and WhisperX transcripts. In doing so, we learn that LLM-based diarization correction models are most effective when correcting transcripts produced using the same ASR as the transcripts they were fine-tuned with, limiting their generalizability. To address this constraint, we build an ensemble model, combining weights from each of the individual models, and demonstrate that such a model has the potential to serve as a reliable and ASR-agnostic diarization correction tool. We hope to make our models available through public APIs for third-party use.
II Methods
Code for all methods can be found in https://github.com/GeorgeEfstathiadis/LLM-Diarize-ASR-Agnostic
II-A Data
@@
II-B Transcription
@@ II-C Data Preprocessing
II-D Model Tuning
II-D1 Model Selection
We used Mistral AI’s Mistral 7b Instruct v0.2 as the base model for all fine-tuned models in this manuscript [37]. Compared to similar open-source large language models, Mistral 7b was chosen because of its size and relative baseline accuracy in diarization correction. The decision-making process behind model selection is presented in Appendix 1.
II-D2 ASR-specific models
Given each ASR tool has a noticeably different transcription style, we hypothesized that a model fine-tuned on one ASR would perform best at correcting diarization for transcripts produced by the same ASR compared to transcripts produced by a different ASR. For this reason, three separate ASR-specific models were fine-tuned using AWS, Azure, and WhisperX transcripts.
@@ II-E Completion parser
II-F Model evaluation
II-F1 Measurement of accuracy
@@ II-F2 Assessing performance
III Results
III-A Transcripts used
@@
The same steps were carried out on the testing set. After preprocessing, we were left with 172 AWS transcripts, 172 Azure transcripts (both AWS and Azure did not have any transcripts removed), and 162 WhisperX transcripts (1 removed due to a single speaker and 9 removed due to repeating word sequences).
III-B Word error rate
When looking at WER across the three ASRs, AWS had a WER of 22.04%, Azure had a WER of 16.99%, and WhisperX had a WER of 22.39%. Azure’s Speech to Text model performed best in terms of transcription accuracy. WhisperX performed the worst.
III-C Baseline diarization accuracy
Diarization accuracy was compared between ASR transcripts and corresponding reference transcripts without any diarization correction. Here, AWS performed best with a deltaCP of 0.93% and deltaSA of 2.5%. Azure was next with a deltaCP of 1.91% and delta SA of 3.06%. WhisperX performed worst, with a deltaCP of 4.46% and deltaSA of 5.77%.
III-D Zero-shot model performance
The performance of the Mistral 7b and Mistral 7x8b models in improving diarization accuracy was measured against baseline, i.e, transcripts with no diarization correction. The results are shown in Table I.
TABLE I: Zero-shot evaluation of base models with no fine-tuning on correction of diarization errors.
@@ III-D Zero-shot model performance
III-E Fine-tuned model performance
The performance of the ASR-specific models and the ensemble model in improving diarization accuracy was measured against baseline. The results are shown in Table II.
TABLE II: Evaluation of fine-tuned models on correction of diarization errors. ASR-specific models performed best on transcripts produced using the same ASR as the transcripts used for fine-tuning. The ensemble model performed best across the board, regardless of the ASR used to produce the transcript.
@@ III-E Fine-tuned model performance
IV Discussion
@@
The ensemble model, with combined weights from each of the individual ASR-specific models, demonstrated better overall performance across ASRs. Evidence suggests the ensemble model may have better awareness of the different types of speaker mislabeling present across ASRs. In the case of correcting diarization errors in the Azure and WhisperX transcripts, the ensemble model performed better than even the ASR-specific models. It may be that different ASR platforms have common error types present in different distributions. A model that is aware of all such types would perform better than an ASR-specific model tuned to identify the more commonly occurring error types in its ASR.
@@
Improved accuracy through the ensemble model was achieved without compromising on inference time. The merging approach did not change the original architecture. Hence, the combined model had the same size as the individual models. We believe an additional advantage of the ensemble model will be that it may generalize better in transcripts from unfamiliar ASRs, as investigated in Appendix 2. This will be useful in real-world applications, where a system may be constrained to a specific cloud platform or have a preference for specific ASR tools. The model may be able to improve diarization accuracy as a post-processing step regardless of the ASR tool used to produce the transcript.
@@
We acknowledge certain limitations, opening up opportunities for future work. First, the performance of diarization correction in an independent dataset remains untested. Different conversational domains may present unique challenges. Second, the project was limited to transcripts in English. This restricts diarization correction to a single language. Future work could test this model on transcripts in different languages and also include additional languages in the dataset used for fine-tuning. Third, our experimentation with varied prompts was not extensive [45]. Given we were building on work from Wang et al. 2024 [24], we used the same prompts utilized in their work. Future work exploring the transferability of the model could include contextual information (e.g., this is a conversation between a doctor and a patient; this is a phone conversation between a salesperson and a customer) [46]. Finally, though our project focused on such models serving as post-processing steps, integrating multimodal data will allow for more robust diarization systems. Combining acoustic information with semantic insights from the LLM will enhance the system’s ability to label speech more effectively [28], especially in complex and noisy environments.
V Conclusion
@@
We investigated the use of fine-tuned LLMs to improve speaker diarization in conversational transcripts with existing speaker labels. Our findings demonstrate that LLMs can markedly improve diarization accuracy when specifically trained to do so. However, if such training is specific to one ASR tool, it may constrain the model’s performance to transcripts only from that ASR tool. An ensemble model with knowledge across ASR tools allows for a generalizable and ASR-agnostic diarization correction tool. We believe such a model can be a useful post-processing step in applications that depend on transcription and diarization of conversations with multiple speakers. We hope to make these models available as public APIs for use by third-party applications.
@@ Appendix A Evaluation of zero-shot pre-trained LLMs
To determine the optimal model for fine-tuning, we conducted an evaluation of various LLMs using a small sample from the test set. A mix of 15 transcriptions was randomly selected, consisting of 5 transcriptions from each ASR (AWS, Azure and WhisperX). After preprocessing and segmenting these transcriptions into manageable prompt-completion pairs, we generated 64 pairs. We focused on small open-source models (fewer than 13 billion parameters), to keep the fine-tuning process computationally feasible and to ensure our results could be made accessible.
@@
TABLE III: Zero-shot evaluation of smaller versions of Llama2 and Mistral base models with no fine-tuning on correction of diarization errors.
@@ Appendix B Generalization to unseen ASR
To assess the generalization capabilities of our models, we transcribed the test set using a fourth ASR, without fine-tuning an ASR-specific model on its training data. We transcribed the test set using the Google Cloud Platform (GCP) transcription service. For the transcription we used the GCP speech to text API (speech_v1p1beta) with the latest_long model enabling speaker diarization capabilities on, setting the maximum speakers to 2 and specifying the language to English. We filtered the testing data by excluding transcriptions with a single speaker (excluded 8) or those with repeated words issue (excluded 0), resulting in 164 transcriptions.
We evaluated all our fine-tuned expert models and the ASR-specific model on the GCP test data, with the results shown in Table IV.
TABLE IV: Evaluation of fine-tuned models on correction of diarization errors in the GCP test set.

### v2

@@ Abstract
Speaker diarization is necessary for interpreting conversations transcribed using automated speech recognition (ASR) tools. Despite significant developments in diarization methods, diarization accuracy remains an issue. Here, we investigate the use of large language models (LLMs) for diarization correction as a post-processing step. LLMs were fine-tuned using the Fisher corpus, a large dataset of transcribed conversations. The ability of the models to improve diarization accuracy in a holdout dataset from the Fisher corpus as well as an independent dataset was measured. We report that fine-tuned LLMs can markedly improve diarization accuracy. However, model performance is constrained to transcripts produced using the same ASR tool as the transcripts used for fine-tuning, limiting generalizability. To address this constraint, an ensemble model was developed by combining weights from three separate models, each fine-tuned using transcripts from a different ASR tool. The ensemble model demonstrated better overall performance than each of the ASR-specific models, suggesting that a generalizable and ASR-agnostic approach may be achievable. We have made the weights of these models publicly available on HuggingFace at https://huggingface.co/bklynhlth.
keywords:
speaker diarization , speech enhancement/separation , multi-speaker , large language model
journal: Computer Speech & Language
[…]
[label1]organization=Brooklyn Health, addressline=11201 USA, city=Brooklyn, state=NY, country=USA
[…]
[label2]organization=Department of Biostatistics, Harvard T. H. Chan School of Public Health, addressline=02115 USA, city=Boston, state=MA, country=USA
[…]
[label3]organization=School of Psychology, University of New South Wales, addressline=2052 Australia, city=Sydney, state=NSW, country=Australia
1 Introduction
@@
Here, we fine-tune LLMs to improve diarization accuracy in ASR transcripts as a post-processing step. To fine-tune our models, we used the English Fisher corpus [31], a large audio dataset containing 1,960 hours of phone conversations, and to test them we used a held out set of the Fisher corpus and the PriMock57 dataset [32] a smaller audio dataset containing mock primary care consultations. We consider the effects of using transcripts from different ASRs, tuning separate models for each AWS, Azure, and WhisperX transcripts. In doing so, we learn that LLM-based diarization correction models are most effective when correcting transcripts produced using the same ASR as the transcripts they were fine-tuned with, limiting their generalizability. To address this constraint, we build an ensemble model, combining weights from each of the individual models, and demonstrate that such a model has the potential to serve as a reliable and ASR-agnostic diarization correction tool. This model has been published on HuggingFace and is available at https://huggingface.co/bklynhlth/WillisDiarize-v1. We additionally released a function using the methods described here for a friendlier user interface on applying diarization correction in OpenWillis [33], a Python library for digital measurement of health, that can be found at https://github.com/bklynhlth/openwillis/wiki/WillisDiarize-with-AWS-v1.0.
2 Methods
Code for all data, training and evaluation methods can be found at https://github.com/GeorgeEfstathiadis/LLM-Diarize-ASR-Agnostic.
2.1 Data
@@
In addition we tested the model on an independent dataset. We used PriMock57 for evaluation [32], a dataset of 57 mock primary consultations between a clinician and a mock patient. The recordings total approximately 9 hours in length, averaging 9 minutes each. There were 7 clinicians and 57 actors portraying the patient involved. Same as in the Fisher corpus, the PriMock57 dataset had corresponding transcripts with sentence-level timestamps. Methods for how these mock consultations were conducted can be found in Korfiatis et al. 2022 [32].
2.2 Transcription
@@ 2.3 Data Preprocessing
2.4 Model Tuning
2.4.1 Model Selection
We used Mistral AI’s Mistral 7b Instruct v0.2 as the base model for all fine-tuned models in this manuscript [39]. Compared to similar open-source large language models, Mistral 7b was chosen because of its size and relative baseline accuracy in diarization correction. The decision-making process behind model selection is presented in A.
2.4.2 ASR-specific models
Given each ASR tool has a noticeably different transcription style, in terms of both words utilized and also diarization errors made, we hypothesized that a model fine-tuned on one ASR would perform best at correcting diarization for transcripts produced by the same ASR compared to transcripts produced by a different ASR. For this reason, three separate ASR-specific models were fine-tuned using AWS, Azure, and WhisperX transcripts.
@@ 2.5 Completion parser
2.6 Model evaluation
2.6.1 Measurement of accuracy
@@ 2.6.2 Assessing performance
3 Results
3.1 Transcripts used
@@
The same steps were carried out on the Fisher testing set and the PriMock57 dataset. After preprocessing, for the Fisher testing set we were left with 172 AWS transcripts, 172 Azure transcripts (both AWS and Azure did not have any transcripts removed), and 162 WhisperX transcripts (1 removed due to a single speaker and 9 removed due to repeating word sequences). While for the PriMock57 dataset we were left with 57 AWS transcripts (0 removed), 57 Azure transcripts (0 removed) and 43 WhisperX transcripts (14 removed due to single or more than 2 speakers).
3.2 Word error rate
We looked at WER across the three ASRs. On the Fisher test set, AWS had a WER of 22.04%, Azure had a WER of 16.99%, and WhisperX had a WER of 22.39%. On the PriMock57 dataset, AWS had a WER of 12.43%, Azure had a WER of 14.28% and WhisperX had a WER of 18.69%. The PriMock57 dataset had considerably better audio quality, given the mock sessions were carefully conducted in an isolated environment, which explains the improved transcription accuracy of the Fisher test set. Azure’s Speech to Text model performed best in terms of transcription accuracy in the Fisher test set, while AWS’s Transcribe performed the best in the PriMock57 dataset. WhisperX performed the worst in both datasets.
3.3 Baseline diarization accuracy
Diarization accuracy was compared between ASR transcripts and corresponding reference transcripts without any diarization correction. In the Fisher dataset, AWS performed best with a deltaCP of 0.93% and deltaSA of 2.5%. Azure was next with a deltaCP of 1.91% and delta SA of 3.06%. WhisperX performed worst, with a deltaCP of 4.46% and deltaSA of 5.77%. On the PriMock57 dataset Azure performed the best with a deltaCP of 0.82% and deltaSA of 1.97%. AWS had a deltaCP of 2.18% and a deltaSA of 3.56%. WhisperX performed worst on the PriMock57 dataset as well with a deltaCP of 4.96% and a deltaSA of 6.56%. The difference in diarization performance between AWS’s Transcribe and Azure’s Speech to Text is indicative of the fact that different ASRs may work better for different use cases.
3.4 Zero-shot model performance
The performance of the Mistral 7b and Mistral 7x8b models in improving diarization accuracy was measured against baseline, i.e, transcripts with no diarization correction. This was done using the Fisher test dataset only. The results are shown in Table 1.
Table 1: Zero-shot evaluation of base models with no fine-tuning on correction of diarization errors in the Fisher test set.
@@ 3.4 Zero-shot model performance
3.5 Fine-tuned model performance
The performance of the ASR-specific models and the ensemble model in improving diarization accuracy was measured against baseline in the Fisher test set. The results are shown in Table 2.
Table 2: Evaluation of fine-tuned models on correction of diarization errors in the Fisher test set. ASR-specific models performed best on transcripts produced using the same ASR as the transcripts used for fine-tuning. The ensemble model performed best across the board, regardless of the ASR used to produce the transcript.
@@ 3.5 Fine-tuned model performance
3.6 Performance on an independent dataset
We also evaluated the fine-tuned models on the PriMock57 dataset. The results are shown in Table 3.
Table 3: Evaluation of fine-tuned models on correction of diarization errors in the PriMock57 dataset. ASR-specific models performed best on transcripts produced using the same ASR as the transcripts used for fine-tuning in most cases. The ensemble model performed best across the board, regardless of the ASR used to produce the transcript.
AWS transcripts Azure transcripts WhisperX transcripts
deltaCP deltaSA deltaCP deltaSA deltaCP deltaSA
Baseline 2.18 3.56 0.82 1.97 4.96 6.56
AWS model 1.79 -18% 3.1 -13% 0.52 -37% 1.63 -17% 4.64 -6% 6.11 -7%
Azure model 1.35 -38% 2.59 -27% 0.46 -44% 1.55 -21% 3.18 -36% 4.53 -31%
WhisperX model 1.33 -39% 2.56 -28% 0.52 -37% 1.62 -18% 2.31 -53% 3.61 -45%
Ensemble model 1.32 -39% 2.56 -28% 0.37 -55% 1.44 -27% 2.68 -46% 3.99 -39%
As seen previously in the Fisher test set, the ASR-specific models from Azure and WhisperX performed the best in the datasets transcribed with the same ASR. The exception was the AWS-specific model which performed the worst in the AWS dataset across ASR-specific models.
The ensemble model achieved the best overall performance in this dataset as well. In the WhisperX transcripts, it had comparable performance to the WhisperX model, which performed the best. However, in the Azure and AWS transcripts, it performed better than all ASR-specific models.
4 Discussion
@@
The performance of the fine-tuned models was influenced by the distribution of speaker diarization errors in the training dataset. When evaluated on an independent dataset, the model’s performance could be negatively affected if the quantity of diarization errors differed significantly from the training dataset, even if transcribed by the same ASR. For instance, the AWS model exhibited the smallest performance improvement on the AWS-transcribed PriMock57 dataset compared to other models. This was likely because AWS-transcribed PriMock57 had a substantially higher baseline rate of speaker diarization errors (deltaCP: 2.18, deltaSA: 3.56) than the AWS-transcribed Fisher test set (deltaCP: 0.93, deltaSA: 2.5). Further analysis revealed that the AWS-specific model became more conservative in the number of speaker changes it implemented, which hindered its performance when evaluated on a dataset with higher error rates.
@@
The ensemble model, with combined weights from each of the individual ASR-specific models, demonstrated better overall performance across ASRs. Evidence suggests the ensemble model may have better adaptability to the different types of speaker mislabeling present across ASRs. Hence, it may be taking a more balanced approach in the amount of speaker corrections it should produce. In the case of correcting diarization errors in the Azure and WhisperX Fisher test transcripts, the ensemble model performed better than the ASR-specific models. It may be that different ASR platforms have common error types present in different distributions. A model that is aware of all such types would perform better than an ASR-specific model tuned to identify the more commonly occurring error types in its ASR.
@@
Improved accuracy through the ensemble model was achieved without compromising on inference time. The merging approach did not change the original architecture. Hence, the combined model had the same size as the individual models. We believe an additional advantage of the ensemble model will be that it may generalize better in transcripts from unfamiliar ASRs, as investigated in B. This will be useful in real-world applications, where a system may be constrained to a specific cloud platform or have a preference for specific ASR tools. The model may be able to improve diarization accuracy as a post-processing step regardless of the ASR tool used to produce the transcript.
We also show that the model is useful on an independent dataset. Notably, this was a clinical dataset, which highlights a major application of such a model i.e. clinical transcriptions. Given the efforts to automate clinical note-taking [47, 48, 49], such a model could improve those tools significantly, particularly if future efforts involve further fine-tuning specific to conversations that consider the use case.
@@
We acknowledge certain limitations, opening up opportunities for future work. First, a more thorough analysis of the diarization errors produced by each ASR, would clarify the need for generalized approaches and help us understand better the types of errors each LLM is more probable to correct. Second, the project was limited to transcripts in English. This restricts diarization correction to a single language. Future work could test this model on transcripts in different languages and also include additional languages in the dataset used for fine-tuning. Third, our experimentation with varied prompts was not extensive [50]. Given we were building on work from Wang et al. 2024 [24], we used the same prompts utilized in their work. Future work exploring the transferability of the model could include contextual information (e.g., this is a conversation between a doctor and a patient; this is a phone conversation between a salesperson and a customer) [51]. Finally, though our project focused on such models serving as post-processing steps, integrating multimodal data will allow for more robust diarization systems. Combining acoustic information with semantic insights from the LLM will enhance the system’s ability to label speech more effectively [28], especially in complex and noisy environments.
5 Conclusion
@@
We investigated the use of fine-tuned LLMs to improve speaker diarization in conversational transcripts with existing speaker labels. Our findings demonstrate that LLMs can markedly improve diarization accuracy when specifically trained to do so. However, if such training is specific to one ASR tool or using a single dataset, it may constrain the model’s performance to transcripts only from that ASR tool and with similar diarization error rates. An ensemble model with knowledge across ASR tools allows for a generalizable and ASR-agnostic diarization correction tool. We believe such a model can be a useful post-processing step in applications that depend on transcription and diarization of conversations with multiple speakers. We have made these models publicly available at HuggingFace at https://huggingface.co/bklynhlth and built a Python function for easier use of the model available at https://github.com/bklynhlth/openwillis/wiki/WillisDiarize-with-AWS-v1.0.
6 Acknowledgements
Though referenced throughout the manuscript, the authors would like to acknowledge the authors of Quan et al. 2024 [24], whose work formed the foundation of this manuscript. Much of the code used during this work, including the TPST algorithm, critical to model fine-tuning and testing, was originally developed by their group and kindly shared with us for replication.
7 Author contributions
Georgios Efstathiadis: Conceptualization, Methodology, Software, Validation, Writing - Original Draft, Visualization Vijay Yadav: Data curation, Software, Investigation, Writing - Review & Editing Anzar Abbas: Writing - Review & Editing, Supervision, Project administration
@@ Appendix A Evaluation of zero-shot pre-trained LLMs
To determine the optimal model for fine-tuning, we conducted an evaluation of various LLMs using a small sample from the test set. A mix of 15 Fisher test transcriptions was randomly selected, consisting of 5 transcriptions from each ASR (AWS, Azure and WhisperX). After preprocessing and segmenting these transcriptions into manageable prompt-completion pairs, we generated 64 pairs. We focused on small open-source models (fewer than 13 billion parameters), to keep the fine-tuning process computationally feasible and to ensure our results could be made accessible.
@@
Table 4: Zero-shot evaluation of smaller versions of Llama2 and Mistral base models with no fine-tuning on correction of diarization errors.
@@ Appendix B Generalization to unseen ASR
To assess the generalization capabilities of our models, we transcribed the test set using a fourth ASR, without fine-tuning an ASR-specific model on its training data. We transcribed the Fisher test set using the Google Cloud Platform (GCP) transcription service. For the transcription we used the GCP speech to text API (speech_v1p1beta) with the latest_long model enabling speaker diarization capabilities on, setting the maximum speakers to 2 and specifying the language to English. We filtered the testing data by excluding transcriptions with a single speaker (excluded 8) or those with repeated words issue (excluded 0), resulting in 164 transcriptions.
We evaluated all our fine-tuned expert models and the ASR-specific model on the GCP test data, with the results shown in Table 5.
Table 5: Evaluation of fine-tuned models on correction of diarization errors in the GCP Fisher test set.